The new project is likely to give a rate of return of 14% before taxes. An investment
company is willing to finance the project through a private placement of 5 crore in the
form of 10% p.a. interest bearing bonds. Apollo’s shares have historically been selling at
P/E of 10. Current earnings are Rs.2.70 and the company is in the 50% tax bracket.
The present Capital structure of the company is:
Long term debt (8%) 1,00,00,000
Common Stock (Rs.2 par value, 1 crore shares outstanding) 2,00,00,000
Retained earnings 17,00,00,000
Total capital 20,00,00,000
before tax marginal cost of the project is 10% (interest on
new loan) and the likely before tax returns from the project are higher than 10% the
company should go ahead. Discuss.
P/E declines to 9, what level of
annual earnings (before interest and taxes) must the new project generate in order to
meet the company’s objective of no change in the value of the stock price.
1
Expert's answer
2020-12-10T13:43:13-0500

IRR > WACC the capital Invested in the investment project will create a return higher than the value of the invested capital. Such a project is attractive for investment

IRR = WACC the Project will not bring any losses or income in the future period and such a project is not attractive

IRR < WACC Such a project will generate negative discounted cash flow in the future

IRR=14

Find WACC:

including shares and bond loans:

"\frac{P}{E}=10"

E=2.7

P=27

"WACC=\frac{E}{K}\times y+\frac{D}{K}\times b\times(1-T)"


K=20 000 000 + 50 000 000= 70 000 000

E=2 000 000

y=10%

D= 50 000 000+1 000 000=51 000 000

"WACC=\frac{2 000 000}{70 000 000}\times10+\frac{1 000 000}{70 000 000}\times 8\times(1-0.5)+\frac{50 000 000}{70 000 000}\times 10\times(1-0.5)=3.91"

"\frac{P}{E}=9"

"WACC=\frac{2 000 000}{70 000 000}\times9+\frac{1 000 000}{70 000 000}\times 8\times(1-0.5)+\frac{50 000 000}{70 000 000}\times 10\times(1-0.5)=3.89"

IRR > WACC

IRR > 3.91